# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ C.5 Müller-Brown potential
[…]
So far, we have demonstrated strong performance of our method in the standard experimental settings considered in the diffusion bridge literature, comparable with the strongest existing method NGDB at a significantly reduced computational cost. We have also observed how taking $\tilde{b}=-b$ in the generalised version of our algorithm can provide improved performance over the standard version in settings where the Jacobian terms become large. To conclude, we consider a harder example in this vein which is chosen to challenge our method and better understand its limitations, inspired by applications in computational chemistry.
The Müller-Brown potential is a classical example of a potential energy surface $U$, and is widely used as a benchmark for studying transition paths of overdamped Langevin dynamics between two metastable states. To examine performance in this setting, we follow the experimental setup used in [15]. The potential used is defined as
$U(x,y)=$ $-200\exp\big(-(x-1)^{2}-10y^{2}\big)$ (39)
$-100\exp\big(-x^{2}-10(y-0.5)^{2}\big)$ (40)
$+15\exp\big(0.7(1+x)^{2}+0.6(x+1)(y-1)+0.7(y-1)^{2}\big),$ (42)
with initial state $x_{0}=[-0.559,1.442]$, terminal state $x_{T}=[0.624,0.028]$, diffusion term $\sigma(t,x)=3{\mathrm{Id}}_{2}$}, and termination time $T=0.05$.
We compare against NGDB, and also to the recent Doob’s Lagrangian method (DL) [15]; this method uses a Gaussian path parameterisation and is specifically designed for transition path sampling, making it a strong baseline for comparison. The steep slopes of the potential make this a challenging setting, so we use $\tilde{b}=-b$ in our method to avoid large Jacobian terms. Indeed, using $\tilde{b}=0$ generally failed to learn the correct dynamics, again reflecting the discussion in Appendix B.4. We used 1000 discretisation steps. For our method and for NGDB, we ran for 4000 training steps, simulated 64 trajectories at each outer step, and used a learning rate of 3e-4; for DL [15] we use the default parameters provided in their codebase. For our method, we also clipped the losses at 1e4, which we found to improve stability.
The results are shown in Table 2, and a visualisation of the obtained trajectories for our method is given in Figure 5. We see that NGDB and our method learn similar values of ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process, though our method has a slightly larger variance across runs. These values are lower than for DL, which is likely because the Gaussian path parameterisation cannot capture the true dynamics as accurately. Their parameterisation does however mean that DL can train significantly faster than our method and NGDB, as it can be trained without requiring simulations. All of the methods learn qualitatively similar trajectories, and have similar maximum energies along the transition paths. We emphasise that we do not aim to achieve state-of-the-art results for transition path sampling (TPS) in this work; this is a field with an extensive literature, with many methods specifically designed to tackle this problem. We instead aim to provide a method for general diffusion bridge problems, and use this example to again highlight the improvements that come from the generalised form of the self-consistency property. Nevertheless, our method obtains good performance in this setting, and extending our method to target TPS problems specifically is a promising direction for future work.
This experiment provides another example showcasing the benefit of taking $\tilde{b}=-b$ in the generalised version of our algorithm, compared to taking $\tilde{b}=0$. This helps to avoid large growth in the Jacobian terms caused by the steep potential. Taking $\tilde{b}=0$ generally caused learning to be unstable and usually fail, other than for some specific choices of $\alpha$-schedule (see the trajectories in Figure 9). In contrast, taking $\tilde{b}=-b$ gave consistent and reliable training.
C.6 $\alpha$-schedule experiments
In Table 3, we perform the same experiments as above, and show the results for the next-step, average, and endpoint $\alpha$-schedules. We use the same hyperparameters as above (aside from adding loss clipping of 10 for the CIR model, for $\tilde{b}=0$ with next-step $\alpha$-schedule). Generally, we find the average schedule to perform strongest so we report those results in the main comparison above; it outperforms the endpoint schedule in all examples, and next-step prediction in the majority. In cases where the Jacobian can lead to large target variance (like in the double-well example, for $\tilde{b}=0$), next-step prediction performs better than average prediction, which is in line with the discussion in Appendix B.4.
[…]
CIR model,
${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$
Double Well,
${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$
Cell Diffusion, ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$$\downarrow$
@@ C.6 $\alpha$-schedule experiments
C.7 Sticking-the-landing adjustments
We now compare performance of our algorithm with and without the STL adjustments discussed in Appendix B.3. We use the same hyperparameters as in the previous experiments, and report the results in Table 4. Including the STL adjustments generally appears to improve performance slightly, with the largest improvements seen in cases where large Jacobians arise (that is, the double-well and Müller-Brown experiments). However, as the STL adjustments require computing the terms $(\nabla u\,\sigma+u\,\nabla\sigma)(t,X_{t}^{u})\cdot\delta B_{t}$, they increase the computational cost of the method and training steps are approximately 1.5 times slower.
@@ C.7 Sticking-the-landing adjustments
CIR model,
${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$
Double Well,
${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$
Cell Diffusion, ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$$\downarrow$
Normal Rare Multimodal

### v2

@@ 4 A General Class of Self-Consistency Losses
CIR model, ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$ Double Well, $v=3.0$, ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$$\downarrow$ Cell Diffusion, ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$$\downarrow$
@@
5 Experiments
To assess the performance of our approach, we consider applying our method CCDB with the two self-consistency properties (SC1) and (4.1), and use the average $\alpha$-schedule (results for other $\alpha$-schedules are reported in Section E.7). We first consider four experimental settings commonly used in the diffusion bridge literature, before finally testing on more challenging examples from computational chemistry. When the ground-truth bridge $\mathbb{P}^{*}$ is available, we report the value of ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$. When it is not, we instead report ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process; amongst processes that bridge between the starting and terminal states, the diffusion bridge is the one that minimises this divergence.
We focus our comparisons on the recent neural-guided method of Yang et al. (2025) (NGDB); this is the most directly comparable method, as it similarly learns a neural control of the drift in an online manner, enabling applications for rare event scenarios in which the alternative neural methods would struggle. Note that we do not aim to necessarily beat the NGDB method, as this method differentiates through the trajectory which provides an advantage. We use NGDB here as a strong baseline and aim to achieve comparable results at a lower computational cost, since our approach does not differentiate through the trajectory. In Table 2, we report runtimes of our algorithm compared to NGDB, and observe an approximate threefold improvement in training speed.
We also include comparisons with the methods score-diffusion bridge (SDB) (Heng et al., 2025), forward-bridge (FB) (Baker et al., 2025), and BEL (Pidstrigach et al., 2025), which learn from uncontrolled simulations. As expected, we find these approaches generally perform well when the $x_{T}$ occurs frequently under the base dynamics, but struggle otherwise, consistent with findings in Yang et al. (2025).
Full details of each experimental setting, additional results, and further discussion of findings are given in Appendix E.
5.1 Ornstein-Uhlenbeck Bridges
Table 2: Comparison of runtimes for our method and NGDB.
Ours NGDB
OU ($\alpha=2.0,d=2$) 230s 800s
CIR 100s 280s
Double Well ($v=3.0$) 180s 520s
Cell Diffusion 60s 170s
Müller-Brown (2dd) 350s 1460s
Müller-Brown (100dd) 1080s 5300s
@@ 5.1 Ornstein-Uhlenbeck Bridges
${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})\downarrow$ Max Energy
@@
We first consider sampling bridges of Ornstein-Uhlenbeck processes ${\mathrm{d}}X_{t}=-\alpha X_{t}{\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}.$ This is a standard setting in the literature for quantitatively evaluating performance (Heng et al., 2025; Baker et al., 2025; Yang et al., 2025), as the true control $u^{*}$ is available in closed-form.
In Figure 3, we compare performance against the NGDB method in various OU bridge settings. We include experimental details and a full discussion of results in Appendix E. In Figure 3(a), we report how the KL divergence relative to the ground truth evolves during training, and see that our method displays improved convergence to the solution in this example. In Figures 3(b) and 3(c), we include plots to illustrate the dependence on dimension and the OU parameter $\alpha$. We found that SDB, FB, and BEL did not learn the bridges successfully in these experiments, which is unsurprising given that they involve rare events under the base dynamics.
5.2 Cox–Ingersoll–Ross model
We now consider computing bridges for the Cox-Ingersoll-Ross (CIR) model (Cox et al., 1985), which is used in mathematical finance to model the evolution of interest rates. The CIR process evolves according to the SDE ${\mathrm{d}}X_{t}=a(b-X_{t}){\mathrm{d}}t+\varepsilon\sqrt{X_{t}}{\mathrm{d}}B_{t}.$} We consider this example as it allows us to verify our method on an SDE with a spatially-dependent diffusion coefficient $\sigma(t,x)=\varepsilon\sqrt{x}$, and also because the ground-truth is again tractable. We find that using (SC1) in our method performs slightly better than using (4.1), which in turn performs comparably with NGDB. The methods SDB, FB and BEL also learn well in this example as the terminal state is well-covered by the base dynamics, though SDB and FB are still outperformed by the methods that learn from controlled dynamics.
5.3 Double-Well Potential
Next, we consider overdamped Langevin dynamics ${\mathrm{d}}X_{t}=-\nabla U(X_{t}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}$} using a classical double-well potential $U(x)=v(x^{2}-1)^{2},$ and aim to learn the bridge from $x_{0}=1$ to $x_{1}=-1$; this setting was previously considered in Pidstrigach et al. (2025). In light of the discussion in Section 4, this presents a potentially more troublesome problem setting for our method, as the concave peak between the two metastable states can lead to large Jacobian terms. We consider setting $v=3.0$ which results in a strong peak between the states, and again report ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$ relative to the ground-truth, which in 1 dimension can be approximated by numerically solving the backwards Kolmogorov equation.
In this example, using (4.1) improves substantially over (SC1), reflecting the discussion in Section 4 regarding large Jacobians. This effect is exacerbated by increased well barrier height $v$; see Section E.3 for a comparison for different barrier heights. The SDB method was significantly outperformed by the approaches that used controlled dynamics, but performed reasonably for wells with lower heights (see Section E.3). We found that FB failed to learn reasonable dynamics in this setting. We also remark that using ‘sticking-the-landing’ adjustments (Roeder et al., 2017; Domingo-Enrich, 2024) provided substantial further improvements for our method in this setting; see Section E.8.
5.4 Cell Diffusion Example
For an example where the reference drift $b$ is not of gradient form, we consider the cell differentiation model of Wang et al. (2011), which has previously been used to provide a qualitative evaluation of diffusion bridge methods in Heng et al. (2025); Baker et al. (2025); Yang et al. (2025). We consider the three settings used in Yang et al. (2025): a normal event, a rare event, and a multi-modal example. There is no ground-truth available in this setting, so we report ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process. We also visualise the obtained trajectories in Figure 6, and observe very similar results for CCDB and NGDB in each case. All methods performed similarly well in the ‘normal’ setting. In the rare and multimodal settings, NGDB appears to perform slightly better than CCDB, but both performed strongly and significantly outperformed the methods that use unconditioned dynamics. SDB and BEL performed reasonably in the rare setting but failed in the multimodal setting, and FB failed to learn the dynamics accurately in both settings.
5.5 Transition Path Sampling
So far, we have demonstrated strong performance of our method in standard experimental settings considered in the diffusion bridge literature, comparable with the strongest existing method NGDB at a significantly reduced computational cost. In the remainder, we apply our method to more challenging examples not previously considered by neural diffusion bridge methods, and aim to assess whether our algorithm remains stable in these more demanding examples.
Transition path sampling (TPS) is a fundamental problem in computational chemistry, in which one aims to simulate the conformational changes of molecules between different metastable states. We focus here on using overdamped Langevin dynamics as we consider elliptic diffusions in this work, and remark that extending our method to the more common underdamped setting is a promising direction for future work (see Section E.6). We emphasise that we do not aim to achieve state-of-the-art performance for TPS; there are many works devoted solely to this problem. Rather, we use this problem as a demanding test case to assess the stability of our method in a more challenging setting.
Müller-Brown Potential
We first consider using the Müller-Brown potential, a classical benchmark for TPS methods. While only two-dimensional, the steep slopes of the potential make this a difficult test case. We compare against NGDB, and also to the recent Doob’s Lagrangian method (DL) (Du et al., 2024); this method uses a Gaussian path parameterisation and is specifically designed for TPS, making it a strong baseline. These approaches impose similar endpoint constraints to our method, making them suitable for comparison. We use the experimental setup of Du et al. (2024); for details, see Section E.5. In CCDB, we use (4.1) to avoid large Jacobian terms; indeed, using (SC1) generally failed to learn the correct dynamics (see Section E.5), consistent with the discussion in Section 4.
The results are shown in Table 3, with trajectories for our method visualised in Figure 4. CCDB and NGDB learn similar values of ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process, though CCDB has a slightly larger variance across runs. These values are lower than for DL, which is likely because the Gaussian path parameterisation cannot capture the true dynamics as accurately. Their parameterisation does however mean that DL can train significantly faster than CCDB and NGDB, as it does not require simulations. All of the methods learn qualitatively similar trajectories, and have similar maximum energies along the transition paths.
In Section E.5, we also consider a 100-dimensional Müller-Brown example, in which we augment the standard potential with quadratic potentials in the additional dimensions. We find that performance remains largely unchanged, demonstrating that problem difficulty is largely dictated by the complexity of the transition rather than dimensionality, and that our method scales well as dimension increases.
Alanine Dipeptide
For an application to a real molecule, we consider transitioning between the $C5$ and C7axC7{ax} conformation states of alanine dipeptide. As we consider overdamped dynamics, our results are not directly comparable with other TPS methods which use underdamped dynamics; nevertheless, the obtained trajectories displayed in Figure 4 resemble those reported in other works. We provide additional details, comparisons with other methods, and discussion of findings in Appendix E.
Choice of Self-Consistency Objective
The experiments shown here demonstrate strong performance of our self-consistency approach for learning diffusion bridges, but also highlight that the best choice of self-consistency property often depends on the behaviour of the Jacobian process. In stable settings without significant growth in the Jacobian terms (the OU, CIR, and cell diffusion examples), the objectives (SC1) and (4.1) perform similarly, whereas in more unstable settings with large Jacobians (double well, Müller-Brown, and Alanine Dipeptide), then (4.1) performs better than (SC1). We provide further experiments to illustrate this effect in Appendix E, which can help to guide practitioners in choosing the best objective for their problem setting.
Figure 4: (Left) Learned trajectories for the Müller–Brown potential, for our method using (4.1). (Middle) Ramachandran plot showing obtained paths for alanine dipeptide. (Right) Energy of an individual transition path for alanine dipeptide.
6 Discussion
Related Work
Diffusion bridges are a well-studied problem with an extended literature spanning several decades (Clark, 1990; Delyon & Hu, 2006; Schauer et al., 2017). Traditional methods have used MCMC (Stuart et al., 2004; Beskos et al., 2008) or SMC-based approaches Lin et al. (2010). Our work contributes to the line of recent methods using neural networks to learn the control of the bridging dynamics (Heng et al., 2025; Baker et al., 2025; Pidstrigach et al., 2025; Yang et al., 2025). An extended discussion of related work is given in Appendix A.
The self-consistency framework provided in this work is also closely related to adjoint matching (Domingo-Enrich et al., 2025) and other recent developments in stochastic optimal control (Domingo-Enrich et al., 2024; Domingo-Enrich, 2024); we discuss these connections in Appendix F.
Limitations
Our approach requires ellipticity assumptions on the base process in order to compute the control terms $u$ in the training objective, which restricts its applicability in certain settings compared to NGDB. The fixed-point nature of our algorithm may be less stable than the KL-based objective used in NGDB, though this could be an inherent trade-off when avoiding differentiation through trajectories. We also require simulations of the current process during training, in contrast to the simulation-free method of Du et al. (2024), though their Gaussian path parameterisation means it does not learn the true dynamics in general.
Conclusion and Future Directions
In this work, we have presented a novel algorithm for learning diffusion bridges based on a self-consistency property of the optimal control $u^{*}$. Our method demonstrates strong performance across a range of diffusion bridge problems, showing competitive performance with the current strongest method at a lower computational cost. Future work could investigate extensions to hypoelliptic diffusions, and target specific applications in computational chemistry for which we report promising findings. In certain settings we observed a significant benefit to using the self-consistency property (4.1); future research could investigate to what extent similar benefits transfer to wider stochastic optimal control problems.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
SH is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning [EP/S023151/1]. JP was supported by the Engineering and Physical Sciences Research Council [EP/Y018273/1].
Appendix
The Appendix is structured as follows. In Appendix A, we present an extended discussion of related work. In Appendix B, we give more details regarding the proposed method and the generalised version discussed in Section 4. In Appendix D we provide the proofs of the results in the paper. In Appendix E, we report experimental details and additional results. In Appendix F, we discuss connections to recent approaches in the stochastic optimal control and generative modelling literature. Finally, in Appendix G we provide the licenses and links to the assets used in the work.
Appendix A Related Work
Diffusion bridge simulation has been the subject of extensive study over recent decades; we highlight a selection of works from the literature here. A common approach is to use simulations from tractable dynamics, and correct using importance-sampling (Papaspiliopoulos & Roberts, 2012) or Metropolis-Hastings steps. Pedersen (1995) use the unconditional dynamics, and Clark (1990); Delyon & Hu (2006) utilise the Brownian bridge to ensure termination at the correct state. Schauer et al. (2017) consider a class of guided diffusions that includes an additional drift term while preserving tractability. We also highlight the works Stuart et al. (2004); Beskos et al. (2008) who consider Langevin-type algorithms on the space of paths, Bierkens et al. (2021) who utilise a piecewise deterministic Monte-Carlo approach, and Lin et al. (2010) who use an SMC-based approach. Often, diffusion bridge approaches are used for inferring the parameters of a discretely-observed diffusion process (Elerian et al., 2001; Roberts & Stramer, 2001; Durham & Gallant, 2002; Stramer & Yan, 2007; Golightly & Wilkinson, 2008; Bladt & Sørensen, 2014; Bladt et al., 2015; van der Meulen & Schauer, 2017).
More recently, neural approximations have been used to learn the control term $u(t,x)$. Once trained, this enables the simulation of samples from the diffusion bridge without the need for MCMC-style procedures. Heng et al. (2025) utilised the expression of the control drift in the time-reversed bridge dynamics as a score function, and trained the drift using score-matching on simulated trajectories of the unconditioned forwards process. The conditioned forwards dynamics can subsequently be learned via a time-reversal. This approach was extended in Baker et al. (2025) to learn the forwards dynamics directly, using trajectories of an adjoint process simulated backwards in time. Pidstrigach et al. (2025) uses a novel characterisation of the control based on Malliavin calculus, which is learned from simulations of the unconditioned process. Such approaches enable direct simulation of the conditioned dynamics using the learned drifts. However, as they rely on simulations of the unconditioned dynamics, they are limited in their ability to learn accurate dynamics for rare events, as these will hardly ever arise in the trajectories used for training. Recent work Yang et al. (2025) aims to address this by using a neural guiding drift in addition to the linear drift from Schauer et al. (2017), and optimising a KL-based objective to encourage towards the true conditioned dynamics. While exhibiting strong performance, the need for backpropagation through simulated trajectories can pose challenges for scalability. It is not clear how to modify their training objective to avoid backpropagation through the trajectories, as simply placing a stop-gradient on the samples from the trajectory will not in general converge to the correct solution. We also highlight the recent work Du et al. (2024), which introduces a variational formulation to learn a Gaussian mixture path parameterisation. Of these methods, the most related to our approach is the recent work of Pidstrigach et al. (2025), as the self-consistency property that we use can also be derived from their Malliavin calculus framework.
Finally, we highlight that the self-consistency framework that we use in this work has many connections to recent advances in the stochastic optimal control literature (Domingo-Enrich et al., 2024, 2025; Domingo-Enrich, 2024), though their focus is on optimal control for rewards with gradient information rather than on diffusion bridges. We provide an extended discussion of these connections in Appendix F.
Appendix B Method Details and Discussion
The presentation of our method in the main text (given in Algorithm 1) is based on the two most natural forms of the self-consistency property, namely (SC1) and (4.1). Below, we include additional implementation considerations regarding our proposed approach, and a detailed description of the general version of the algorithm using the generalised self-consistency property discussed in Section 4.
B.1 Choice of Neural Parameterisation
Recall from the discussion in Section 3 that we use the neural parameterisation to guide towards the terminal state, and to enforce the gradient-form property. Aside from these considerations, there is considerable freedom in the choice of neural parameterisation, though it is clearly beneficial to direct the samples towards the endpoint to avoid large jumps in the final step. As discussed in Section 3, in cases with a gradient-form base drift $b$ we use the parameterisation
$f_{\theta}(t,X_{t})=\frac{x_{T}-X_{t}}{T-t}+\sigma(t,X_{t})\,\eta_{\theta}(t,X_{t}),$ (16)
for a neurally-parameterised vector field $\eta$. Recalling the result in Theorem 3.1, we only need to enforce the gradient property at a single point in time to give a theoretically-grounded algorithm. In cases with gradient-form $b$ and diffusion term $\sigma$ independent of $x$ (which covers a significant proportion of use cases), this is ensured by our construction where we directly jump to the terminal state in the final step; this enforces that the control $u(T-\delta t,\cdot)$ is of gradient-form. In practice, rotational terms may accrue if the optimisation is not performed perfectly; in many settings we found improved performance by enforcing the gradient property throughout time by setting $\eta_{\theta}=\nabla\psi_{\theta}$ for scalar-valued network $\psi_{\theta}$.
For base drifts $b$ not of gradient-form, we include the drift $b$ in the parameterisation,
$f_{\theta}(t,X_{t})=b(t,X_{t})+\frac{x_{T}-X_{t}}{T-t}+\sigma(t,X_{t})~\eta_{\theta}(t,X_{t}).$
As above, we can enforce the gradient-form property directly throughout time by setting $\eta_{\theta}=\nabla\psi_{\theta}$. If one wishes to use a more general neural parameterisation $\eta$, note that for time close to $T$ the samples are close to the terminal point $x_{T}$ and thus the drift $b$ can be well-approximated by its linearisation about $x_{T}$. Thus, jumping to the terminal state in the final step approximately enforces $u(T-\delta t,\cdot)$ to be of gradient-form similarly to above (and we note that we found such design decisions to have minimal effect in practice).
Certain choices of spatially-dependent diffusion coefficients $\sigma$ may require further modifications, and the appropriate adjustments can again be made to ensure the gradient-form conditions are satisfied in these cases too. There are also further possibilities for the parameterisation, such as incorporating the linearised guiding drift used in Mider et al. (2021); Yang et al. (2025).
B.2 Solving for the training targets
Recall that we compute the training targets $\mathcal{U}_{s}:=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t$ by solving backwards along the simulated trajectories. To do so, we can write a recursive formula for the target $\mathcal{U}_{s}$ in terms of the next-step target $\mathcal{U}_{s+\delta t}$. We include the full calculations below. Note that here (and in similar computations throughout the Appendix) we write $X_{t}$ for notational simplicity, and recall that these computations are performed using the controlled dynamics $X_{t}^{u_{\theta}}$.
$\mathcal{U}_{s}=\tfrac{1}{A_{s}}\int_{s}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t$ (17)
$=\tfrac{1}{A_{s}}\int_{s+\delta t}^{T}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{1}{A_{s}}\int_{s}^{s+\delta t}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t$ (18)
$=\tfrac{1}{A_{s}}\int_{s+\delta t}^{T}\alpha_{t}(J_{t|s+\delta t}J_{s+\delta t|s})^{\top}\,u(t,X_{t}){\mathrm{d}}t+\tfrac{1}{A_{s}}\int_{s}^{s+\delta t}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t$ (19)
$=\tfrac{A_{s+\delta t}}{A_{s}}J_{s+\delta t|s}^{\top}\,\mathcal{U}_{s+\delta t}+\tfrac{1}{A_{s}}\int_{s}^{s+\delta t}\alpha_{t}J_{t|s}^{\top}\,u(t,X_{t}){\mathrm{d}}t$ (20)
$\approx J_{s+\delta t|s}^{\top}\,\big[\tfrac{A_{s+\delta t}}{A_{s}}\mathcal{U}_{s+\delta t}+(\delta t)\tfrac{\alpha_{s+\delta t}}{A_{s}}u(s+\delta t,X_{s+\delta t})\big],$ (21)
where in the final line we approximate the second integral using its value at the larger timestep. By relabelling $\lambda_{s}=\tfrac{A_{s+\delta t}}{A_{s}}$, this can be written as
$\mathcal{U}_{s}=J_{s+\delta t|s}^{\top}\,\big[\lambda_{s}\mathcal{U}_{s+\delta t}+(1-\lambda_{s})u_{s+\delta t}(X_{s+\delta t})\big].$ (22)
This is intuitive; each target $\mathcal{U}_{s}$ is constructed as a weighted combination of the ‘running’ next-step target $\mathcal{U}_{s+\delta t}$ and the next-step control $u(s+\delta t,X_{s+\delta t})$, and is then transformed by the single-step Jacobian $J_{s+\delta t|s}\approx{\mathrm{Id}}+\nabla b(s,X_{s})(\delta t)+\nabla\sigma(s,X_{s})\cdot\delta B_{s}$. The choice of $\alpha$-schedule therefore determines how ‘far into the future’ the algorithm looks to enforce the self-consistency property. Following Pidstrigach et al. (2025), we consider three different choices of $\alpha$-schedule: next-step, average, and endpoint prediction. For clarity, we explicitly write out the recursions for these three $\alpha$-schedules below.
  • Next-step prediction: This choice of schedule chooses to only enforce the self-consistency property at consecutive timesteps; that is, it tries to enforce $u(s,X_{s}^{u})=\mathbb{E}[J_{s+\delta t|s}^{\top}\,u(s+\delta t,X_{s+\delta t}^{u})|X_{s}^{u}]$. This corresponds to taking $\lambda_{s}=0$ in the above expression (22), so the targets are calculated as
    $\mathcal{U}_{s}=J_{s+\delta t|s}^{\top}\,u(s+\delta t,X_{s+\delta t}^{u}).$ (23)